OpenStudy (anonymous):

Can anyone state the fundamental theorem of calculus part 2?

OpenStudy (dumbcow):

\[\int\limits_{a}^{b} f(x) dx = F(b) - Fa)\] where \[F'(x) = f(x)\]

OpenStudy (anonymous):

Thank you

OpenStudy (dumbcow):

yw